OpenStudy (anonymous):

Solve . Answer 5 1 4 2

OpenStudy (anonymous):



OpenStudy (anonymous):

@jennychan12

OpenStudy (jennychan12):

the 4 isn't in the radical?

OpenStudy (anonymous):

I know but it just solve

OpenStudy (anonymous):

I think it 2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Correct me if I'm wrong..

OpenStudy (anonymous):

\[\sqrt x^2 = 2^2\]\[x=4\]

OpenStudy (jennychan12):

\[\sqrt{x-4} = 2 \] or \[\sqrt{x} - 4 = 2\] ? it makes a difference what the answer is.

OpenStudy (anonymous):

So x equal 4?

OpenStudy (anonymous):

is that what I typed

OpenStudy (anonymous):

then yes

OpenStudy (anonymous):

Yeah thanks But I wanna be sure...

OpenStudy (jennychan12):

if it's \[\sqrt{x-4} = 2\] then the answer is 8 it it's \[\sqrt{x}-4=2\] then x = 36

OpenStudy (anonymous):

it's equal to -2 btw